How many kilograms of water must be evaporated from 40 kilograms of 10% solution to produce a 25% salt solution?
Answer by rwm(914)

40/10=x/4
4=x/4
16=x
check
40*.1=.25*16
4=4
ok
the 25% solution will be 16 kilo
you must evaporate 24 kilos